Circular No. 7044

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

     URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1998 U5

Frank Shelly reports that one of the recent LINEAR discoveries is clearly a comet. The following precise CCD observations are available:

1998 UT α_{2000} δ_{2000} m_{1} Observer
Oct. 30.37456 6^{h}53^{m}34.^{s}08 +28^{o}57′57′′.5 14.0 Blythe
30.38514 6 53 32.32 +28 58 58.5  _{”}
30.39590 6 53 29.51 +28 59 50.6  _{”}
30.40675 6 53 27.55 +29 00 40.7  _{”}
30.41759 6 53 24.83 +29 01 36.5  _{”}
30.98976 6 51 15.61 +29 48 38.5  Manca
30.99595 6 51 14.09 +29 49 11.4  _{”}
31.07056 6 50 56.18 +29 55 29.9 14.3 Marechal
31.07340 6 50 55.38 +29 55 44.2 14.5  _{”}
31.07608 6 50 54.75 +29 55 57.8 14.3  _{”}
31.07862 6 50 54.19 +29 56 10.5 14.1  _{”}
31.08992 6 50 51.39 +29 57 08.4 14.4  _{”}
31.11426 6 50 45.56 +29 59 13.0 13.9  _{”}
31.11666 6 50 44.95 +29 59 24.2 14.0  _{”}
31.20729 6 50 22.74 +30 07 07.7 14.6  _{”}
31.20854 6 50 22.36 +30 07 14.7 14.0  _{”}
31.30463 6 49 59.38 +30 15 30.8 12.3 Kowalski
31.31233 6 49 57.58 +30 16 10.1 12.5  _{”}
31.31852 6 49 55.94 +30 16 42.3 12.5  _{”}

M. Blythe, F. Shelly, and M. Bezpalko (Lincoln Laboratory ETS). Measur-

ers J. Stuart, H. Viggh, R. Sayer. 1.0-m f/2.15 reflector. F. Manca and M. Cavagna (Sormano). 0.5-m reflector. Diffuse coma and

2′ westward tail. F. Marechal (Observatoire de Dax). 0.20-m reflector. Tail ∼ 2′ in p.a. 260^{o}. R. A. Kowalski (Zephyrhills). 0.18-m f/9 reflector.

COMET C/1998 K5 (LINEAR)

Visual m_{1} estimates: Sept. 20.05 UT, 12.2 (M. Meyer, Frauenstein, Germany, 0.25-m reflector); Oct. 2.42, 12.9 (A. Hale, Cloudcroft, NM, 0.41m reflector); 19.81, 13.0 (A. Pearce, Nedlands, W. Australia, 0.41-m reflector); 25.79, 13.5 (W. Hasubick, Buchloe, Germany, 0.44-m reflector); 27.01, 14.2 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector).

1998 October 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green